OpenStudy (anonymous):

Using the completing the square method find the solutions of g(x)? G(x)=x^2-8x+12 please help me i dont get this at all

OpenStudy (dumbcow):

x^2 - 8x + 12 = 0 x^2 -8x = -12 take half of 8 and square it, 4^2 =16 .... add to both sides x^2 -8x + 16 = -12 + 16 (x-4)^2 = 4 x - 4 = +-2 x = 2,6

OpenStudy (anonymous):

@dumbcow thank you so much i really appreciate it

OpenStudy (dumbcow):

yw ... use that same process every time you solve using completing the square

OpenStudy (anonymous):

@dumbcow ok can you help me with 2 things quick

OpenStudy (dumbcow):

sure

OpenStudy (anonymous):

@dumbcow 1. explain if completing the square is a good method for solving when the discriminant is negative? 2. use h(x)=x^2+6x+12 as an example to justify your claim algebraicly and thanks again

OpenStudy (dumbcow):

well there are 2 methods for solving quadratics.... completing the square or quadratic formula In either case if the discriminant is negative then no real solutions exist, solutions will be imaginary.

OpenStudy (anonymous):

@dumbcow so you cant use the method

OpenStudy (dumbcow):

you can use either method, i dont know what they mean by asking if its a good method

OpenStudy (anonymous):

@dumbcow ok what about #2

OpenStudy (dumbcow):

solve it the same way... what is half of 6 squared?

OpenStudy (anonymous):

@dumbcow would it be 18 cause 6^2 is 36 half of that is 18

OpenStudy (dumbcow):

no just like in the first problem "x" coefficient was 8 ... we took half of 8, then squared it .... 4^2 = 16 here the "x" coefficient is 6 , half of 6, then square it ..... 3^2 = 9

OpenStudy (anonymous):

@dumbcow then you just add it to both sides

OpenStudy (dumbcow):

yes x^2 +6x + 9 = -12 + 9

OpenStudy (anonymous):

@dumbcow ok now this is where im stuck now what do i do

OpenStudy (dumbcow):

look at the first example i did. you have to factor left side, in general its always (x+b/2)^2 where "b" is middle coefficient

OpenStudy (anonymous):

@dumbcow im sorry im confused on what you do next

OpenStudy (dumbcow):

do you remember factoring?

OpenStudy (anonymous):

@dumbcow if i have to guess is it (x+6/2)^2 yes i do do you factor he number

OpenStudy (dumbcow):

right, but what is 6/2 ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

(x+3)^2 = -3

OpenStudy (anonymous):

o

OpenStudy (anonymous):

@dumbcow now what do we do

OpenStudy (anonymous):

we have (x+3)^2=-3

OpenStudy (anonymous):

would it be x=-3

OpenStudy (dumbcow):

lol you didn;t understand my steps in the first problem huh ok we had (x-4)^2 = 4 then x-4 = +-2 what operation cancels out a square?

OpenStudy (anonymous):

divide?

OpenStudy (dumbcow):

nope

OpenStudy (anonymous):

how would u do it

OpenStudy (dumbcow):

solve this x^2 = 16 x = ?

OpenStudy (anonymous):

x=4

OpenStudy (dumbcow):

how did you get that? what operation did you use on the 16

OpenStudy (anonymous):

square root

OpenStudy (dumbcow):

there you go...to cancel out (x+3)^2 you have to take square root

OpenStudy (anonymous):

so it would be x+3= im stuck now @dumbcow

OpenStudy (anonymous):

SQRT of -3

OpenStudy (dumbcow):

yep which is \[\sqrt{3} i\]

OpenStudy (dumbcow):

sorry it should be plus or minus as well

OpenStudy (anonymous):

so it equals \[x +3=\sqrt{3i}\]

OpenStudy (anonymous):

+- before the SQRT

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

oh ok great thank you @dumbcow

OpenStudy (dumbcow):

\[x = \pm i \sqrt{3} - 3\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@dumbcow have a good nighht thank you